# Establishing Linear Surrogate Regret Bounds for Convex Smooth Losses via Convolutional Fenchel–Young Losses

### v2

@@ Abstract
Surrogate regret bounds, also known as excess risk bounds, bridge the gap between the convergence rates of surrogate and target losses, with linear bounds favorable for their lossless regret transfer. While convex smooth surrogate losses are appealing in particular due to the efficient estimation and optimization, the existence of a trade-off between the smoothness and linear regret bound has been believed in the community. That being said, the better optimization and estimation properties of convex smooth surrogate losses may inevitably deteriorate after undergoing the regret transfer onto a target loss. We overcome this dilemma for arbitrary discrete target losses by constructing a convex smooth surrogate loss, which entails a linear surrogate regret bound composed with a tailored prediction link. The construction is based on Fenchel–Young losses generated by the convolutional negentropy, which are equivalent to the infimal convolution of a generalized negentropy and the target Bayes risk. Consequently, the infimal convolution enables us to derive a smooth loss while maintaining the surrogate regret bound linear. We additionally benefit from the infimal convolution to have a consistent estimator of the underlying class probability. Our results are overall a novel demonstration of how convex analysis penetrates into optimization and statistical efficiency in risk minimization.
@@ 1 Introduction
The risk of a machine learning model is often measured by the expectation of a target loss $\ell$ that quantifies the error between the model prediction $t$ and a natural label $y$. However, minimizing such a target risk over a dataset is often computationally hard because a target prediction problem is usually discretely structured, including multiclass, multilabel, top-$k$ prediction problems. For this reason, a tractable surrogate risk induced by a surrogate loss $L(\bm{\theta},y)$ serves as an essential proxy, where $\bm{\theta}$ is a score vector that often takes value in $\mathbb{R}^{\varrho}$.
An ideal surrogate loss should be convex, smooth, and calibrated toward a given target prediction problem. Convexity and smoothness (or Lipschitz continuous gradient) of a loss function have been fundamental properties both in optimization and statistical estimation—indeed, classical optimization theory reveals that convex and smooth functions can be efficiently optimized with first-order methods [47], which is prevalent in modern machine learning. In addition, several studies have demonstrated that the convexity and smoothness can further enhance fast rates in the risk estimation [63, 71]. Meanwhile, calibration is regarded as a minimal requirement on surrogate losses, ensuring that the surrogate risk minimization leads to the target risk minimization [6, 64].111 Readers should distinguish the calibration property in this context from the calibration of probabilistic predictions [24]. To establish a relationship between surrogate and target risks, surrogate regret bounds play a crucial role. Therein the regret (or the suboptimality) of a target loss is controlled by that of a surrogate loss through a non-decreasing rate function $\psi:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$. A surrogate regret bound can be informally written as follows: for any score vector $\bm{\theta}\in\mathbb{R}^{\varrho}$ and a class distribution $\bm{\eta}$,
@@
where $\varphi$ is a prediction link that converts a score vector $\bm{\theta}$ into a target prediction $\varphi(\bm{\theta})$. If the regret rate function is linear $\psi(r)=\mathcal{O}(r)$, which is the best possible (data-independent) regret rate,222 A super-linear regret bound is possible if we allow a surrogate loss to depend on a data distribution [70], which is not considered here. then the optimization and estimation errors of the surrogate regret are optimally translated to the target regret. For example, under the binary classification target loss, Bartlett et al. [6] demonstrated that the linear regret convergence is possible when the hinge loss is used as a surrogate loss; additionally, symmetric losses [15] and polyhedral losses [25] yield the linear regret convergence. Unfortunately, these loss functions lack either convexity or smoothness, which deteriorates the optimization and estimation errors of the surrogate regret, even if they enjoy linear regret bounds. By contrast, a square-root regret rate function $\psi(r)=\mathcal{O}(\sqrt{r})$ is commonly encountered when it comes to convex smooth surrogate losses, such as the logistic, exponential, and squared losses [6, 49, 25]. These loss functions enjoy better optimization and estimation properties, yet suffer from larger target regrets due to the suboptimal regret rate. Therefore, it has been open for a long time to develop a convex smooth surrogate loss without sacrificing the regret convergence rate.
Notwithstanding, the previous literature in this line has implied that such an ideal surrogate loss may be inconceivable. Mahdavi et al. [33] considered this question with an interpolated loss between the hinge (non-smooth) and logistic (smooth) losses and showed that we inevitably face the trade-off between generalization and optimization unless strong distributional assumptions are imposed. Further, Frongillo and Waggoner [25] proved that locally smooth and strongly convex losses must suffer from a square-root regret convergence rate at least. Ramaswamy et al. [53] blame convex smooth losses for redundantly modeling continuous class distributions, which is unnecessary if our goal is merely to solve discrete target problems.
In this paper, we demonstrate that this seemingly impossible trade-off can be overcome for every discrete target loss $\ell$. Specifically, we build a convex smooth surrogate loss built upon the framework of Fenchel–Young losses [10]—a framework to generate a loss function from a generalized negentropy and its conjugate. In a nutshell, our main results are summarized as follows:
@@
For any discrete target loss $\ell$, there exist a convex smooth surrogate loss $L$ (defined over a score $\bm{\theta}\in\mathbb{R}^{\varrho}$) and prediction link $\varphi$ such that the surrogate regret bound (1) holds with some linear rate $\psi(r)=\mathcal{O}(r)$.
@@
In addition, we improve the multiplicative term in the initial linear surrogate regret bound in Section 3.4 to make the bound tighter, which exploits the low-rank structure of a target loss $\ell$. As a by-product of the loss smoothness, we can provide a Fisher-consistent probability estimator of the underlying probability in Section 3.5. Finally, Section 4 instantiates the framework of the convolutional Fenchel–Young loss for the multiclass classification problem, highlighting the efficient computation of the prediction link.
@@ 1.1 Related Works
A discrete prediction problem aims to predict $t$ that minimizes a discrete target loss $\ell(t,y)$ over the class distribution, and the study of convex surrogates for $\ell$ is of significant interest. Extensive research has been conducted on surrogate losses for specific discrete tasks, including but not limited to classification [6, 74, 65, 67, 61, 46, 43], top-$k$ [69, 66], and multilabel learning [26, 73, 37, 32, 44].
@@
In contrast to ad-hoc approaches, recent studies have advanced the principled design of calibrated convex surrogate losses for general discrete prediction problems, without imposing restrictions on the target losses. In Ramaswamy and Agarwal [51] and Ramaswamy et al. [54], surrogate losses based on the squared loss and error correcting output codes are proposed, respectively, which embed the structure of a target problem into a surrogate loss and are shown to be calibrated with the corresponding target losses. The design of calibrated surrogate losses has also been extensively studied in the context of structured prediction [19, 52, 17, 18, 50, 48, 49], where the structural/low-rank properties of the target losses are exploited to construct more efficient surrogates. As opposed to the smooth losses utilized in the works above, polyhedral losses [21, 22, 23] has attracted attention recently, which provides a systematic framework for constructing efficiently calibrated yet non-smooth convex losses based on a discrete target loss.
Previous studies on surrogate regret bounds.
@@
Surrogate regret bounds have been well studied for margin losses under binary and multiclass classification [74, 6, 65, 60, 39, 62], where the target loss is the 0-1 loss. Similarly, proper losses [14, 27, 56, 2, 29, 45] have been shown to provide surrogate regret bounds w.r.t. the $L_{1}$ distance between the estimated and true class probabilities [55, 4, 5], which facilitates analyses for downstream tasks. For structured prediction, surrogate regret bounds are sometimes called comparison inequalities [17, 50, 48, 8], which are typically of square-root type. Notably, Frongillo and Waggoner [25] demonstrated that polyhedral losses exhibit linear surrogate regret bounds for general discrete prediction problems, covering various piecewise linear non-smooth losses [69, 53] as special cases. However, they lack the smoothness. To the contrary, Mao et al. [36] and Mao et al. [35] propose smooth losses with linear regret rates but lacking convexity.
@@
Whereas a growing line of research has focused on $\mathcal{H}$-consistency to analyze how the restriction to a function class $\mathcal{H}$ affects consistency [31, 72, 3], we adopt surrogate regrets over all measurable score functions because the analysis is often more transparent and it is reasonable to suppose that the function class is sufficiently expressive under the overparametrization regime. The extension to $\mathcal{H}$-consistency is rather straightforward by integrating the minimizability gap [35, 38, 37, 40, 34].
@@ 2 Preliminaries
Let $[d]\coloneqq\{1,2,\cdots,d\}$ and [[]]delimited-[]delimited-[][\![$\cdot$]\!][ [ ⋅ ] ] is the Iverson bracket. The $p$-norm is denoted by $\|\cdot\|_{p}$, which we assume to be the $2$-norm unless otherwise noted. Let $\overline{\mathbb{R}}\coloneqq\mathbb{R}\cup\{\infty\}$ be the extended real-line and $\Delta^{d}\coloneqq\{\bm{\eta}\in\mathbb{R}^{d}_{\geq 0}:\|\bm{\eta}\|_{1}=1\}$ the $d$-simplex. For a set $\mathcal{S}\subseteq\mathbb{R}^{d}$, $\mathrm{int}(\mathcal{S})$ and $\mathrm{relint}(\mathcal{S})$ are its interior and relative interior, respectively, and $\mathrm{conv}(S)$ is its convex hull. The indicator function is denoted by $\mathbb{I}_{\mathcal{S}}:\mathbb{R}^{d}\to\{0,+\infty\}$, where $\mathbb{I}_{\mathcal{S}}(\bm{\theta})=0$ if $\bm{\theta}\in\mathcal{S}$ and +$\infty$+ ∞ otherwise. For a function $f:\mathbb{R}^{d}\to\overline{\mathbb{R}}$, $\mathrm{dom}(\Omega)\coloneqq\{\bm{\theta}\in\mathbb{R}^{d}:\Omega(\bm{\theta})<+\infty\}$ is its effective domain. A function $f$ is extended to be set-valued with slight abuse of notation by $f(\mathcal{S})\coloneqq\{f(s):s\in\mathcal{S}\}$. The Fenchel conjugate of $\Omega$ is $\Omega^{*}(\bm{\theta})\coloneqq\sup_{\bm{p}\in\mathrm{dom}(\Omega)}\bm{\theta}^{\top}\bm{p}-\Omega(\bm{p})$. The canonical basis of the Euclidean space is denoted by $\{\bm{e}_{i}\}$, where the dimensionality depends on the contexts.
@@ 2.1 Discrete Prediction Problems and Target Losses
In this section, we review the target losses that define discrete prediction problems, and further introduce their equivalent lower-dimensional decompositions.
Discrete prediction problem:
Let $\mathcal{Y}=[K]$ be the finite class space. The class distribution on $\mathcal{Y}$ is $\bm{\eta}\in\Delta^{K}$ such that class $Y=y$ with probability $\eta_{y}$. A discrete prediction problem aims to find a target prediction $t$ from the finite prediction space $\widehat{\mathcal{Y}}\coloneqq[N]$ for each $\bm{\eta}$ that minimizes the expectation of a discrete target loss $\ell:\widehat{\mathcal{Y}}\times\mathcal{Y}\to\mathbb{R}$ over $y\sim\bm{\eta}$. This expectation is called the risk w.r.t. the target loss $\ell$: $R_{\ell}(t,\bm{\eta})\coloneqq\mathbb{E}_{y\sim\bm{\eta}}[\ell(t,y)]=\langle\bm{\eta},\bm{\ell}(t)\rangle$, where $\bm{\ell}(t)\in\mathbb{R}^{K}$ is the loss vector such that $\bm{\ell}(t)_{y}=\ell(t,y)$. The Bayes risk of $\ell$ at $\bm{\eta}$ is $\underline{R}_{\ell}(\bm{\eta})\coloneqq\min_{t\in\widehat{\mathcal{Y}}}R_{\ell}(t,\bm{\eta})$, which is a concave function of $\bm{\eta}$. The suboptimality of a prediction $t$ w.r.t. the target loss $\ell$ over a class distribution $\bm{\eta}$ is characterized by the target regret, which is the gap between its risk and the Bayes risk.
@@
An equivalent but more efficient decomposition of (3) called Structure Encoding Loss Functions (SELF) is introduced in Ciliberto et al. [17], which allows lower-dimensional encoding of class labels. It is often leveraged by follow-up works to induce efficient loss functions, and is further generalized by the Affine Decomposition in Blondel [8, (12)]. In this work, we will also represent discrete target losses in this decomposed form.
@@
where $\bm{\rho}:\mathcal{Y}\to\mathbb{R}^{\varrho}$ is a label encoding function that maps discrete labels into the $\varrho$-dimensional Euclidean space, $\bm{\ell}^{\bm{\rho}}:\widehat{\mathcal{Y}}\to\mathbb{R}^{\varrho}$ is the corresponding  loss encoding function, and the term $c:\mathcal{Y}\to\mathbb{R}$ is the remainder independent of prediction $t$.
This loss decomposition contains the Affine Decomposition [8, (12)], which generalizes SELF [17]. Any discrete target loss enjoys this decomposition since it can be trivially decomposed by $\bm{\rho}(y)=\bm{e}_{y}$, $\bm{\ell}^{\bm{\rho}}(t)=\bm{\ell}(t)$, and $c=0$ with $\varrho=K$, which immediately recovers (3). With a properly chosen $\bm{\rho}$, the encoded cardinality $\varrho$ can often be greatly smaller than $K$, particularly in the context of structured prediction. For example, consider multilabel classification, where $\mathcal{Y}=\widehat{\mathcal{Y}}=[K]$, $K=2^{\varrho}$, and each $y\in[K]$ corresponds to a unique binary vector $\bm{\nu}(y)\in\{0,1\}^{\varrho}$ representing a set of binary labels. Suppose our target loss is the Hamming loss $\ell_{H}(t,y)\coloneqq\bm{1}^{\top}(\bm{\nu}(t)+\bm{\nu}(y))-2\langle\bm{\nu}(t),\bm{\nu}(y)\rangle$, which is the Hamming distance between $\bm{\nu}(t)$ and $\bm{\nu}(y)$. This entails the decomposition $\bm{\rho}(y)=\bm{1}-2\bm{\nu}(y)$, $\bm{\ell}^{\bm{\rho}}(t)=\bm{\nu}(t)$, and $c(y)=\bm{1}^{\top}\bm{\nu}(y)$. Here, $\varrho=\ln_{2}K$ significantly reduces the dimensionality from the cardinality of $\mathcal{Y}$. We refer reader to Blondel [8, Appendix A] for more examples of discrete prediction problems. The cardinality $\varrho$ of a label encoding function $\bm{\rho}$ is closely related to surrogate regret bounds later in Section 3.4.

### v3

@@ Abstract
Surrogate regret bounds, also known as excess risk bounds, bridge the gap between the convergence rates of surrogate and target losses. The regret transfer is lossless if the surrogate regret bound is linear. While convex smooth surrogate losses are appealing in particular due to the efficient estimation and optimization, the existence of a trade-off between the loss smoothness and linear regret bound has been believed in the community. Under this scenario, the better optimization and estimation properties of convex smooth surrogate losses may inevitably deteriorate after undergoing the regret transfer onto a target loss. We overcome this dilemma for arbitrary discrete target losses by constructing a convex smooth surrogate loss, which entails a linear surrogate regret bound composed with a tailored prediction link. The construction is based on Fenchel–Young losses generated by the convolutional negentropy, which are equivalent to the infimal convolution of a generalized negentropy and the target Bayes risk. Consequently, the infimal convolution enables us to derive a smooth loss while maintaining the surrogate regret bound linear. We additionally benefit from the infimal convolution to have a consistent estimator of the underlying class probability. Our results are overall a novel demonstration of how convex analysis penetrates into optimization and statistical efficiency in risk minimization.
@@ 1 Introduction
The risk of a machine learning model is often measured by the expectation of a target loss $\ell$ that quantifies the error between the model prediction $t$ and a natural label $y$. However, minimizing the target risk over a dataset is often computationally hard because a target prediction problem is usually discretely structured, including multiclass, multilabel, top-$k$ prediction problems. For this reason, a tractable surrogate risk induced by a surrogate loss $L(\bm{\theta},y)$ serves as an essential proxy with a score $\bm{\theta}\in\mathbb{R}^{d}$. The resulting surrogate optimization is no longer discretely constrained.
An ideal surrogate loss should be convex, smooth (or entailing a Lipschitz continuous gradient), and calibrated toward a given target prediction problem. Convexity and smoothness have been fundamental both in optimization and statistical estimation—indeed, classical optimization theory reveals that convex smooth functions can be optimized with first-order methods more efficiently than non-smooth functions [yuri, SAG]. In addition, several studies have demonstrated that the convexity and smoothness can further enhance fast rates in the risk estimation [smooth1, smooth2]. Meanwhile, calibration is regarded as a minimal requirement on surrogate losses, ensuring that the surrogate risk minimization leads to the target risk minimization [06_calibration, steinwart2007compare]. [^1] Readers should distinguish this calibration property from calibrated prediction [foster_calibration]. To establish a relationship between surrogate and target risks, surrogate regret bounds play a crucial role. Therein the regret (or the suboptimality) of a target loss is controlled by that of a surrogate loss through a non-decreasing rate function $\psi:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$. A surrogate regret bound can be informally written as follows: for any score vector $\bm{\theta}\in\mathbb{R}^{d}$ and a class distribution $\bm{\eta}$,
@@
where $\varphi$ is a prediction link that converts a score vector $\bm{\theta}$ into a target prediction $\varphi(\bm{\theta})$. If the regret rate function is linear $\psi(r)=\mathcal{O}(r)$, which is the best possible (data-independent) regret rate, [^2] A super-linear bound is possible with distribution-dependent losses [zhang2021rates], which we do not consider here. then the optimization and estimation errors of the surrogate loss are optimally translated to the target loss. For example, under the binary classification target loss, 06_calibration demonstrated that the linear regret rate is possible with the hinge loss. Later, symmetric losses [Nontawat] and polyhedral losses [poly_regret] were shown to yield the linear regret rate. Unfortunately, these loss functions lack either convexity or smoothness, which deteriorates the optimization and estimation errors of the surrogate regret, even if they enjoy linear regret bounds. By contrast, a square-root regret rate $\psi(r)=\mathcal{O}(\sqrt{r})$ is common for convex smooth surrogate losses, such as the logistic, exponential, and squared losses [06_calibration, Self5, poly_regret]. These losses typically enjoy better optimization and estimation properties, yet suffer from larger target regrets due to the suboptimal regret rate. Therefore, it has been open to develop a convex smooth surrogate loss without sacrificing the linear regret rate.
Notwithstanding, the previous literature in this line has implied that such an ideal surrogate loss may be inconceivable. mahdavi considered this question with an interpolated loss between the hinge (non-smooth) and logistic (smooth) losses and showed that we inevitably face the trade-off between generalization and optimization unless strong distributional assumptions are imposed. Further, poly_regret proved that locally smooth and strongly convex losses must suffer from a square-root regret rate at least. rama_BEP blame convex smooth losses for redundantly modeling continuous class distributions, which is unnecessary if our goal is merely to solve discrete target problems.
In this paper, we demonstrate that this seemingly impossible trade-off can be overcome for arbitrary discrete target losses. Specifically, we build a convex smooth surrogate loss built upon the framework of Fenchel–Young losses [FY]—a framework to generate a loss function from a generalized negentropy and its conjugate. In a nutshell, our main results are summarized as follows:
@@
For any discrete target loss $\ell$, there exist a convex smooth surrogate loss $L$ (defined over a score $\bm{\theta}\in\mathbb{R}^{d}$) and prediction link $\varphi$ such that the surrogate regret bound (1) holds with some linear rate $\psi(r)=\mathcal{O}(r)$.
@@
In addition, we improve the multiplicative term in the initial linear surrogate regret bound in Section 3.4 to make the bound tighter, which exploits the low-rank structure of a target loss $\ell$. As a by-product of the loss smoothness, we can provide a Fisher-consistent probability estimator of the underlying probability in Section 3.5. Finally, Section 4 instantiates the framework of the convolutional Fenchel–Young loss for the multiclass classification problem, highlighting the efficient computation of the prediction link. More examples of target prediction problems, such as classification with rejection and multilabel ranking, are available in Appendix D.
@@ 1.1 Related Works
A discrete prediction problem aims to predict $t$ that minimizes a discrete target loss $\ell(t,y)$ over the class distribution, and the study of convex surrogates for $\ell$ is of significant interest. Extensive research has been conducted on surrogate losses for specific discrete tasks, including but not limited to classification [06_calibration, 04_multi, 07_multi, yutong, scott_asy, hari_multi, menon_bin, f_diver], top-$k$ [koyejo_topk, top_k_poly], and multilabel learning [multilabel1, multilabel2, mao_multi, menon_bip, menon_multi].
@@
In contrast to ad-hoc approaches, recent studies have advanced the principled design of calibrated convex surrogate losses for general discrete prediction problems, without imposing restrictions on the target losses. In rama_cvxdim and rama_ecoc, surrogate losses based on the squared loss and error correcting output codes are proposed, respectively, which embed the structure of a target problem into a surrogate loss and are shown to be calibrated with the corresponding target losses. The design of calibrated surrogate losses has also been extensively studied in the context of structured prediction [cortes2016structured, Self0, Self1, Self2, Self_osokin, Self4, Self5, cao2024consistent], where the structural/low-rank properties of target losses are exploited to construct more efficient surrogates. As opposed to the smooth losses utilized in the works above, polyhedral losses [poly1, poly_dim, total_poly] have attracted attention recently, which provides a systematic framework for efficiently constructing calibrated yet non-smooth convex losses based on a discrete target loss.
Surrogate regret bounds.
@@
Surrogate regret bounds have been well studied for margin losses under binary and multiclass classification [04_multi, 06_calibration, 07_multi, cs_regret, maobin, jessie], where the target loss is the 0-1 loss. Similarly, proper losses [proper1, proper2, proper3, regret_proper4, pmlr-v45-Kotlowski15, menon_auc] have been shown to provide surrogate regret bounds w.r.t. the $L_{1}$ distance between the estimated and true class probabilities [regret_proper1, regret_proper2, regret_proper3], which facilitates analyses for downstream tasks. For structured prediction, surrogate regret bounds are sometimes called comparison inequalities [Self1, Self_osokin, Self4, Self3], which are typically of square-root type. Notably, poly_regret demonstrated that polyhedral losses exhibit linear surrogate regret bounds for general discrete prediction problems, covering various piecewise linear non-smooth losses [koyejo_topk, rama_BEP] as special cases. However, they lack the smoothness. To the contrary, mao_st and mao_cs propose smooth losses with linear regret rates but lacking convexity.
@@
Whereas a growing line of research has focused on $\mathcal{H}$-consistency to analyze how the restriction to a hypothesis space $\mathcal{H}$ affects consistency [hc1, hc2, hc3], we implicitly suppose that the hypothesis space is all measurable functions because the analysis is often more transparent and it is reasonable to suppose that the hypothesis space is sufficiently expressive under the overparametrization regime. The extension to $\mathcal{H}$-consistency is rather straightforward by integrating the minimizability gap [mao_cs, mao_reject, mao_multi, H_cons1, H_cons2]. [^3] This condition can be relaxed to the well-specified hypothesis space, under which the hypothesis space is required to contain at least one population risk minimizer, rather than the entire space of measurable functions.
@@ 2 Preliminaries
Let $d]\coloneqq\{1,2,\cdots,d\}$ and $\![\cdot$]\!] is the Iverson bracket. The $p$-norm is denoted by $\|\cdot\|_{p}$, which we assume to be the $2$-norm unless otherwise noted. Let $\overline{\mathbb{R}}\coloneqq\mathbb{R}\cup\{\infty\}$ be the extended real-line and $\Delta^{d}\coloneqq\{\bm{\eta}\in\mathbb{R}^{d}_{\geq 0}:\|\bm{\eta}\|_{1}=1\}$ the $d$-simplex. For a set $\mathcal{S}\subseteq\mathbb{R}^{d}$, $\mathrm{int}(\mathcal{S})$ and $\mathrm{relint}(\mathcal{S})$ are its interior and relative interior, respectively, and $\mathrm{conv}(\mathcal{S})$ is its convex hull. The indicator function is denoted by $\mathbb{I}_{\mathcal{S}}:\mathbb{R}^{d}\to\{0,+\infty\}$, where $\mathbb{I}_{\mathcal{S}}(\bm{\theta})=0$ if $\bm{\theta}\in\mathcal{S}$ and $\infty$ otherwise. For a function $f:\mathbb{R}^{d}\to\overline{\mathbb{R}}$, $\mathrm{dom}(f)\coloneqq\{\bm{\theta}\in\mathbb{R}^{d}:f(\bm{\theta})<+\infty\}$ is its effective domain. A function $f$ is extended to be set-valued with slight abuse of notation by $f(\mathcal{S})\coloneqq\{f(s):s\in\mathcal{S}\}$. The Fenchel conjugate of $\Omega$ is $\Omega^{*}(\bm{\theta})\coloneqq\sup_{\bm{p}\in\mathrm{dom}(\Omega)}\{\bm{\theta}^{\top}\bm{p}-\Omega(\bm{p})\}$. The identity matrix is $I$. The canonical basis of the Euclidean space is denoted by $\{\bm{e}_{i}\}$, where the dimensionality depends on the contexts.
@@ 2.1 Discrete Prediction Problems and Target Losses
Let $\mathcal{Y}=[K]$ be the finite class space. The class distribution on $\mathcal{Y}$ is $\bm{\eta}\in\Delta^{K}$ such that class $Y=y$ has probability $\eta_{y}$. A discrete prediction problem aims to find a target prediction $t$ from the finite prediction space $\widehat{\mathcal{Y}}\coloneqq[N]$ for each $\bm{\eta}$ by minimizing a discrete target loss $\ell:\widehat{\mathcal{Y}}\times\mathcal{Y}\to\mathbb{R}$ over $y\sim\bm{\eta}$. The averaged target loss is called the target risk: $R_{\ell}(t,\bm{\eta})\coloneqq\mathbb{E}_{y\sim\bm{\eta}}[\ell(t,y)]=\langle\bm{\eta},\bm{\ell}(t)\rangle$, where $\bm{\ell}(t)\in\mathbb{R}^{K}$ is the loss vector such that $\bm{\ell}(t)_{y}=\ell(t,y)$ for each $y\in\mathcal{Y}$. The Bayes risk of $\ell$ at $\bm{\eta}$ is $\underline{R}_{\ell}(\bm{\eta})\coloneqq\min_{t\in\widehat{\mathcal{Y}}}R_{\ell}(t,\bm{\eta})$, which is a concave function of $\bm{\eta}$. The suboptimality of a prediction $t$ w.r.t. the target loss $\ell$ over a class distribution $\bm{\eta}$ is characterized by the target regret, which is the gap between its risk and the Bayes risk.
@@
An equivalent but more efficient decomposition of (3), known as Structure Encoding Loss Functions (SELF), was introduced in Self1, which allows lower-dimensional label encodings. It has been widely used to construct efficient loss functions, and further generalized via Affine Decomposition in Self3. We also adopt a general form to represent discrete target losses.
@@
where $\bm{\rho}:\mathcal{Y}\to\mathbb{R}^{d}$ is a label encoding function that maps discrete labels into the $d$-dimensional Euclidean space, $\bm{\ell}^{\bm{\rho}}:\widehat{\mathcal{Y}}\to\mathbb{R}^{d}$ is the corresponding loss encoding function, and $c:\mathcal{Y}\to\mathbb{R}$ is the remainder independent of prediction $t$.
This loss decomposition contains the Affine Decomposition [Self3, (12)]. Any discrete target loss admits such a decomposition via the trivial choice $\bm{\rho}(y)=\bm{e}_{y}$, $\bm{\ell}^{\bm{\rho}}(t)=\bm{\ell}(t)$, and $c=0$ with $d=K$, which immediately recovers (3). With a properly chosen $\bm{\rho}$, the encoded cardinality $d$ can often be greatly smaller than $K$, particularly in the context of structured prediction. For example, consider multilabel classification, where $\mathcal{Y}=\widehat{\mathcal{Y}}=[K]$, $K=2^{d}$, and each $y\in[K]$ corresponds to a unique binary vector $\bm{\nu}(y)\in\{0,1\}^{d}$ representing a set of binary labels. Suppose our target loss is the Hamming loss $\ell_{H}(t,y)\coloneqq\bm{1}^{\top}(\bm{\nu}(t)+\bm{\nu}(y))-2\langle\bm{\nu}(t),\bm{\nu}(y)\rangle$, which is the Hamming distance between $\bm{\nu}(t)$ and $\bm{\nu}(y)$. This entails the decomposition $\bm{\rho}(y)=\bm{1}-2\bm{\nu}(y)$, $\bm{\ell}^{\bm{\rho}}(t)=\bm{\nu}(t)$, and $c(y)=\bm{1}^{\top}\bm{\nu}(y)$. Here, $d=\log_{2}K$ significantly reduces the dimensionality from the cardinality of $\mathcal{Y}$. We refer reader to Self3 and Appendix D for more examples of discrete prediction problems. The cardinality $d$ of label encoding $\bm{\rho}$ is closely related to surrogate regret bounds later in Section 3.4.